# Equivariant scaling asymptotics for Poisson and Szegő kernels on Grauert tube boundaries

### v1

@@ 4.1 General arguments for the scaling asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}$
::=:= $\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}}\cdot F_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}\right];$
@@ 4.3 Proof of Theorem 25
Let us apply (4.1), (4.1), and (4.1) with $x_{1,\lambda}$ and $x_{2,\lambda}$ swapped and $\chi$ replaced by $\chi_{-}$. Then $t_{1}$ is replaced with $-t_{1}$, and $B$ with $B^{-1}$; hence $P$ gets replaced with $P^{*}$ in (114) (see the proof of Proposition 4.17 of [F]). Therefore, $\left|\sigma_{-t_{1},0}^{\tau}(x_{2})\right|=\left|\sigma_{t_{1},0}^{\tau}(x_{1})\right|$. More precisely, by Lemma 78,
@@
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
@@
where $\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})$ and $\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ are as in Definition 23 (here $g_{l}=\kappa_{l}\,h_{1}$).
@@
$-\left\langle\mathbf{u},\frac{1}{\tau}J_{0}B^{-1}J_{0}\mathbf{v}\right\rangle+\imath\,\left\langle\mathbf{u},\frac{1}{\tau}\,J_{0}\,B^{-1}\,\mathbf{v}\right\rangle-\frac{1}{2}\,\langle\mathbf{u},R\,\mathbf{u}\rangle,$
@@
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot\mathrm{det}(R)^{-\frac{1}{2}}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
@@
$\sim$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi,\lambda}(x_{1},x_{2})\cdot\sum_{l=1}^{r_{x_{1}}}e^{-\frac{1}{2}\,\langle\mathbf{V}(\tau),D_{l}\,\mathbf{V}(\tau)\rangle}$
@@
$\delta^{\prime}_{\lambda}$ ::=:= $\max\left\{\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},Z^{\tau}\right),\mathrm{dist}_{X^{\tau}}\left(x_{2,\lambda},Z^{\tau}\right)\right\}$
@@
$\delta^{\prime\prime}_{\lambda}$ ::=:= $\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},x_{2,\lambda}^{G\times\mathrm{supp}(\chi)}\right)$
@@ 5 Scaling asymptotics for $P_{k,\nu,\lambda}$
    $\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{O}(X^{\tau}),$
    which is a Fourier integral operator with complex phase of degree $-(d-1)/4$;
@@
    The composition $U_{\mathbb{C}}(t+2\,\imath\,\tau):=\mathfrak{P}^{\tau}\circ U(t)\circ{\mathfrak{P}^{\tau}}^{*}$ is a Fourier integral operator with complex phase of degree $-(d-1)/2$ on $X^{\tau}$, whose distributional kernel admits the spectral description
@@
  8. 8.
@@
    then on $X^{\tau}$ there exist a smoothly varying pseudodifferential operator $Q^{\tau}_{t}$ of degree $-(d-1)/2$ and a smoothly varying operator $R^{\tau}_{t}$ with $\mathcal{C}^{\infty}$ kernel such that
@@
    $\gamma^{\tau}_{t,0}(x)=\tau^{\frac{d-1}{2}}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}\cdot\langle\sigma_{J}^{(x)},\sigma_{J_{t}}^{(x)}\rangle^{-1}$ (151)
@@
  11. 11.
    The principal symbol of $D^{\tau}_{\sqrt{\rho}}$ along $\Sigma^{\tau}$ (or, equivalently, of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$) is
@@
The arguments in the proof of Theorem 7 apply, except that the leading term of the amplitude has been multiplied by a factor $\tau^{(d-1)/2}\,(u\,\tau)^{-(d-1)/2}$ and the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ in (151) replaces $e^{\imath\,\theta^{\tau}_{t}(x)}$. In view of the rescaling $u\mapsto\lambda\,u$, this entails a further factor $\lambda^{-(d-1)/2}$. Furthermore, at the critical point (101) of the phase the product $u\,\tau=1$

### v3

@@ 4.1 General arguments for the scaling asymptotics of $\Pi^{\tau}_{\chi,\nu,\lambda}$
$:=$ $\int_{\mathfrak{g}}\,\mathrm{d}\xi\,\int_{\mathbb{R}^{2d-2}}\,\mathrm{d}\mathbf{u}\,\left[e^{\mathcal{S}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}}\cdot F_{k}(\theta_{1},\mathbf{v}_{1},\theta_{2},\mathbf{v}_{2};\mathbf{u},\xi)_{l}\right];$
@@ 4.3 Proof of Theorem 25
Let us apply (4.1), (4.1), and (4.1) with $x_{1,\lambda}$ and $x_{2,\lambda}$ swapped and $\chi$ replaced by $\chi_{-}$. Then $t_{1}$ is replaced with $t_{1}$, and $B$ with $B^{-1}$; hence $P$ gets replaced with $P^{*}$ in (114) (see the proof of Proposition 4.17 of [F]). Therefore, $\left|\sigma_{-t_{1},0}^{\tau}(x_{2})\right|=\left|\sigma_{t_{1},0}^{\tau}(x_{1})\right|$. More precisely, by Lemma 78,
@@
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
@@
where $\mathcal{F}_{\chi}(x_{1},x_{2})$ and $\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$ are as in Definition 23 (here $g_{l}=\kappa_{l}\,h_{1}$).
@@
$\left\langle\mathbf{u},\frac{1}{\tau}J_{0}B^{-1}J_{0}\mathbf{v}\right\rangle+\imath\,\left\langle\mathbf{u},\frac{1}{\tau}\,J_{0}\,B^{-1}\,\mathbf{v}\right\rangle-\frac{1}{2}\,\langle\mathbf{u},R\,\mathbf{u}\rangle,$
@@
$\asymp$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot\mathrm{det}(R)^{-\frac{1}{2}}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\mathcal{B}_{\nu}(x_{1},x_{2})_{l}$
@@
$\sim$ $\frac{e^{-\imath\,\lambda\,t_{1}}}{\sqrt{2\,\pi}}\,\left(\frac{\lambda}{2\,\pi\,\tau}\right)^{d-1-d_{G}/2}\cdot e^{\imath\,\theta^{\tau}_{t_{1}}(x_{1})}\cdot\mathcal{F}_{\chi}(x_{1},x_{2})\cdot\sum_{l=1}^{r_{x_{1}}}e^{-\frac{1}{2}\,\langle\mathbf{V}(\tau),D_{l}\,\mathbf{V}(\tau)\rangle}$
@@
$\delta^{\prime}_{\lambda}$ $:=$ $\max\left\{\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},Z^{\tau}\right),\mathrm{dist}_{X^{\tau}}\left(x_{2,\lambda},Z^{\tau}\right)\right\}$
@@
$\delta^{\prime\prime}_{\lambda}$ $:=$ $\mathrm{dist}_{X^{\tau}}\left(x_{1,\lambda},x_{2,\lambda}^{G\times\mathrm{supp}(\chi)}\right)$
@@ 5 Scaling asymptotics for $P_{k,\nu,\lambda}$
    $\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{O}(X^{\tau}),$ (148)
    which is a Fourier integral operator with complex phase of degree (d1)/4-(d-1)/4;
@@
    The composition $U_{\mathbb{C}}(t+2\,\imath\,\tau):=\mathfrak{P}^{\tau}\circ U(t)\circ{\mathfrak{P}^{\tau}}^{*}$ is a Fourier integral operator with complex phase of degree (d1)/2-(d-1)/2 on $X^{\tau}$, whose distributional kernel admits the spectral description
@@
For the following computations, we need the description (due to Zelditch) of of the complexified Poisson-wave operators $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ as a ‘dynamical Toeplitz operator’. The latter relies in turn on the study of the composition
$\mathfrak{Q}^{\tau}:={\mathfrak{P}^{\tau}}^{*}\circ\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{C}^{\infty}(M),$
where $\mathfrak{P}^{\tau}$ is as in (148). As explained in §3 of [Z-2007], $\mathfrak{Q}^{\tau}$ is an elliptic pseudodifferential operator on $M$, of degree (d1)/2-(d-1)/2, and its principal symbol $\sigma(\mathfrak{Q}^{\tau})$ appears in the description of $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ in terms of dynamical Toeplitz operators (see [Z-2012], [Z-2014], [Z-2020]).
Since $\mathfrak{P}^{\tau}$$^{*}$ depends on the choice of volume form on $X^{\tau}$, so does $\mathfrak{Q}^{\tau}$. In this section, we review the computation of $\sigma(\mathfrak{Q}^{\tau})$, which was carried out by Zelditch in §3 of [Z-2007], in light of the choices that we have adopted (see also the discussion in [P-2024]). We follow the general heuristic strategy in [Z-2007]: first consider the Euclidean case, and then reduce the general case to the latter.
Thus we first assume that $M=\mathbb{R}^{d}$ with the standard metric, and $X^{\tau}\cong\mathbb{R}^{d}\times S^{d-1}_{\tau}$ (where $S^{d-1}_{\tau}=\tau\,S^{d-1}$ is the sphere of radius $\tau$ centered at the origin).
In the real domain, the wave kernel at time $t$ for $\mathbb{R}^{d}$ with the standard metric is
$E_{t}(x,y):=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,(t\,\|\xi\|+\langle\xi,x-y\rangle)}\,\mathrm{d}\xi.$ (150)
This may be analytically continued to the complex domain in $t$ and $x$ by replacing $t$ with $t+\imath\,\tau$ for $\tau>0$ and $x$ with $\zeta=x+\imath\,p\in\mathbb{C}^{d}$. In particular, for $t=0$ we obtain the kernel
$\mathfrak{E}^{\tau}(\zeta,y)=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{-\tau\,\|\xi\|-\langle\xi,p\rangle+\imath\,\langle\xi,x-y\rangle}\,\mathrm{d}\xi,$ (151)
which is absolutely convergent and holomorphic in $\zeta$ on the locus where $\|p\|<\tau$ (which plays the role of $M^{\tau}$). The distributional kernel of $\mathfrak{P^{\tau}}$ is given by the restriction of (151) to the locus where $\|p\|=\tau$ (which plays the role of $X^{\tau}$).
With respect to the given volume forms, $(\mathfrak{P}^{\tau})^{*}$ is represented by the integral kernel
$(\mathfrak{P}^{\tau})^{*}(y,\zeta)=\overline{\mathfrak{P}^{\tau}(\zeta,y)}=\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{-\tau\,\|\xi\|-\langle\xi,p\rangle-\imath\,\langle\xi,x-y\rangle}\,\mathrm{d}\xi,$
where $\zeta=x+\imath\,p$.
Therefore $\mathfrak{Q}^{\tau}$ is represented by the operator kernel
$\mathfrak{Q}^{\tau}(x,y)=\int_{X^{\tau}}(\mathfrak{P}^{\tau})^{*}(x,\zeta)\,\,\mathfrak{P}^{\tau}(\zeta,y)\,\mathrm{d}V_{X^{\tau}}(\zeta).$ (152)
Let us write $\zeta=x^{\prime}+\imath\,\tau\,\omega$, where $x^{\prime}\in\mathbb{R}^{d}$ and $\omega\in S^{d-1}$. Then $\mathrm{d}V_{X^{\tau}}(x^{\prime})=\tau^{d-1}\,\mathrm{d}x^{\prime}\,\mathrm{d}\omega$. We obtain
$\mathfrak{Q}^{\tau}(x,y)$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\left[(\mathfrak{P}^{\tau})^{*}(x,\zeta)\,\,\mathfrak{P}^{\tau}(\zeta,y)\,\right]$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{1}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{2}$
$\left[e^{-\tau\,(\|\xi_{1}\|+\|\xi_{2}\|)-\tau\,\langle\xi_{1}+\xi_{2},\omega\rangle+\imath\,\langle\xi_{2},x^{\prime}-y\rangle-\imath\,\langle\xi_{1},x^{\prime}-x\rangle}\,\right]$
$=$ $\frac{\tau^{d-1}}{(2\,\pi)^{2\,d}}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}x^{\prime}\,\int_{S^{d-1}}\mathrm{d}\omega\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{1}\,\int_{\mathbb{R}^{d}}\,\mathrm{d}\xi_{2}$
$\left[e^{-\tau\,(\|\xi_{1}\|+\|\xi_{2}\|)-\tau\,\langle\xi_{1}+\xi_{2},\omega\rangle+\imath\,\langle\xi_{2}-\xi_{1},x^{\prime}\rangle+\imath\,\langle\xi_{1},x\rangle-\imath\,\langle\xi_{2},y\rangle}\,\right].$
Using the distributional identity
$\frac{1}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,\langle\xi_{2}-\xi_{1},x^{\prime}\rangle}\,\mathrm{d}x^{\prime}=\delta(\xi_{2}-\xi_{1}),$
we can rewrite (5) as
$\mathfrak{Q}^{\tau}(x,y)=\frac{\tau^{d-1}}{(2\,\pi)^{d}}\,\int_{\mathbb{R}^{d}}\,e^{\imath\,\langle\xi,x-y\rangle}\left[\int_{S^{d-1}}e^{-2\,\tau\,\|\xi\|-2\,\tau\,\langle\xi,\omega\rangle}\,\mathrm{d}\omega\right]\mathrm{d}\xi.$ (154)
We aim to evaluate the inner integral asymptotically for $\xi\rightarrow\infty$. To this end, we set $\xi=\lambda\,\eta$, where $\lambda>0$ and $\eta\in S^{d-1}$, and let $\lambda\rightarrow+\infty$. We obtain
$\int_{S^{d-1}}e^{-2\,\tau\,\|\xi\|-2\,\tau\,\langle\xi,\omega\rangle}\,\mathrm{d}\omega$ $=$ $\int_{S^{d-1}}e^{\imath\lambda\cdot\Psi_{\eta}^{\tau}(\omega)}\,\mathrm{d}\omega,$ (155)
where
$\Psi_{\eta}^{\tau}(\omega):=\imath\,2\,\tau\,\big{(}1+\langle\eta,\omega\rangle\big{)}.$
Thus $\Psi_{\eta}^{\tau}$ is purely imaginary, and $\Im\left(\Psi_{\eta}^{\tau}\right)\geq 0$. Furthermore, $\Im\left(\Psi_{\eta}^{\tau}\right)$ vanishes only at $\omega=-\eta$. Hence, without altering the asymptotics for $\lambda\rightarrow+\infty$, we may replace integration over $S^{d-1}$ by integration over an arbitrarily small open neighbourhood $S_{\eta}\subset S^{d-1}$ of $\eta$.
Furthermore, given the form of $\Psi_{\eta}^{\tau}:S^{d-1}\rightarrow\mathbb{R}$, there is no loss of generality in assuming that $\eta$ is the last vector of the standard basis of $\mathbb{R}^{d}$, that is, $\eta=\begin{pmatrix}\mathbf{0}&1\end{pmatrix}^{\dagger}$. Thus any $\omega\in S_{\eta}$ may be written $\omega=\begin{pmatrix}\mathbf{u}&-\sqrt{1-\|\mathbf{u}\|^{2}}\end{pmatrix}^{\dagger}$, where $\mathbf{u}\in\mathbb{R}^{d-1}$ ranges in a small neighbourhood of the origin. Then $\mathrm{d}\omega=\mathcal{V}(\mathbf{u})\,\mathrm{d}\mathbf{u}$ on $S_{\eta}$ where $\mathcal{V}(\mathbf{0})=1$. For $\mathbf{u}\sim\mathbf{0}$,
$\Psi_{\eta}^{\tau}(\omega)=\imath\,2\,\tau\,\left(1-\sqrt{1-\|\mathbf{u}\|^{2}}\right)=\imath\,\tau\,\left(\|\mathbf{u}\|^{2}+R_{3}(\mathbf{u})\right).$ (156)
Thus there is a unique critical point $\mathbf{u}=\mathbf{0}$ (that is, $\omega=-\eta$), with Hessian matrix $H(\Psi_{\eta}^{\tau})=2\,\imath\,\tau\,I_{d-1}$.
Thus
$\sqrt{\det\left(\frac{\lambda\,H(\Psi_{\eta}^{\tau})}{2\pi\,\imath}\right)}=\left(\frac{\lambda\tau}{\pi}\right)^{\frac{d-1}{2}}.$
Applying the Lemma of stationary phase, we obtain for (155) an aymptotic expansion in descending powers of $\lambda=\|\xi\|$, with leading order term $(\pi/\lambda\tau)^{(d-1)/2}$. In view of (154) the principal symbol of $\mathfrak{Q}^{\tau}$ is therefore
$\sigma(\mathfrak{Q}^{\tau})=\tau^{d-1}\,\left(\frac{\pi}{\lambda\,\tau}\right)^{\frac{d-1}{2}}=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}.$ (157)
Before considering the general case, let us premise a remark considering scalar rescalings of a metric. Suppose $X=X^{\tau}$ and consider the operator $\mathfrak{P}^{\tau}:\mathcal{C}^{\infty}(M)\rightarrow\mathcal{O}(X)$. To emphasize the role of the metric, let us write $\mathfrak{P}^{\tau}=\mathfrak{P}^{\tau}_{\kappa}$, $X=X^{\tau}_{\kappa}$.
When $\kappa$ is replaced by $\kappa^{\lambda}:=\lambda^{2}\,\kappa$ for some $\lambda\in\mathbb{R}_{+}$, we have $X=X^{\lambda\,\tau}_{\kappa^{\lambda}}$. If $\Delta_{\kappa}$ and $\Delta_{\kappa^{\lambda}}$ are the Laplacians for $\kappa$ and $\kappa^{\lambda}$, respectively, then $\Delta_{\kappa^{\lambda}}=\lambda^{-2}\,\Delta_{\kappa}$. Hence,
$e^{-\tau\,\sqrt{\Delta_{\kappa}}}=e^{-\lambda\,\tau\,\sqrt{\Delta_{\kappa^{\lambda}}}}.$
Thus it makes sense to denote $\mathfrak{P}^{\tau}_{\kappa}$ as $\mathfrak{P}^{X}$, without reference to a specific rescaling of $\kappa$.
On the other hand, the rescaling affects the adjoint operator, since it modifies the volume form on $X$. Let $\mathrm{vol}_{X,\kappa}$ and let $\mathrm{vol}_{X,\kappa^{\lambda}}$ be the volume forms on $X$ viewed as $X^{\tau}_{\kappa}$ and $X^{\lambda\,\tau}_{\kappa^{\lambda}}$. Similarly, let $\mathrm{vol}_{M,\kappa}$ and $\mathrm{vol}_{M,\kappa^{\lambda}}$ be the Riemannian volume forms on $M$ associated to $\kappa$ and $\kappa^{\lambda}$, respectively. Then
$\mathrm{vol}_{M,\kappa^{\lambda}}=\lambda^{d}\,\mathrm{vol}_{M,\kappa},\quad\mathrm{vol}_{X,\kappa^{\lambda}}=\lambda^{2d-1}\,\mathrm{vol}_{X,\kappa}.$ (158)
Let $({\mathfrak{P}}^{X})^{*}_{\kappa}$ and $({\mathfrak{P}}^{X})^{*}_{\kappa^{\lambda}}$ be the adoints of $\mathfrak{P}^{X}$ with respect to $\kappa$ and $\kappa^{\lambda}$ (that is, using the pairs of volume forms $(\mathrm{vol}_{M,\kappa},\mathrm{vol}_{X,\kappa})$, $(\mathrm{vol}_{M,\kappa^{\lambda}},\mathrm{vol}_{X,\kappa^{\lambda}})$, respectively). One concludes from (158) that
$({\mathfrak{P}}^{X})^{*}_{\kappa^{\lambda}}=\lambda^{d-1}\,({\mathfrak{P}}^{X})^{*}_{\kappa}.$
Since $X=X^{\tau}_{\kappa}=X^{\lambda\,\tau}_{\kappa^{\lambda}}$, we obtain
$\mathfrak{Q}^{\lambda\,\tau}_{\kappa^{\lambda}}=\lambda^{d-1}\,\mathfrak{Q}^{\tau}_{\kappa};$ (159)
therefore the same relation holds between the respective principal symbols. Since $\mathfrak{Q}^{\tau}_{\kappa}$ has degree (d1)/2-(d-1)/2, we conclude that its principal symbol has the form
$\sigma(\mathfrak{Q}^{\tau}_{\kappa})(m,\xi)=c_{m}(\xi)\,\left(\frac{\tau\,\pi}{\|\xi\|}\right)^{\frac{d-1}{2}},$ (160)
where $c_{m}(\xi)$ is homogeneous of degree 0 in $\xi$.
Let us consider a general real-analytic $(M,\kappa)$ and fix $m\in M$. Let us choose a real-analytic coordinate chart $\varphi:B\rightarrow U$ centered at $m$ that is isometric at the origin. Thus, $B\subseteq\mathbb{R}^{d}$ is an open neighborhood of the origin (say, an open ball centered at $\mathbf{0}$), $U$ an open neighbourhood of $m$ in $M$, and $\varphi$ a real-analytic diffeomorphism such that $\mathrm{d}_{\mathbf{0}}\varphi$ is a linear isometry between $(\mathbb{R}^{d},g_{st})$ (where $g_{st}$ is the standard Euclidean product) and $(T_{m}M,\kappa_{x})$. The pull-back metric $\varphi^{*}(\kappa)$ admits a convergent power series expansion
$\varphi^{*}(\kappa)_{\mathbf{x}}=g_{st}+\sum_{|I|\geq 1}x^{I}\,g_{I}\qquad(\mathbf{x}\in B),$
where the $g_{I}$’s are fixed symmetric 2-tensors. Here $x^{I}=x_{1}^{i_{1}}\cdots x_{d}^{i_{d}}$ if $\mathbf{x}=\begin{pmatrix}x_{1}&\ldots&x_{d}\end{pmatrix}^{\dagger}$.
For some sufficiently small $\epsilon>0$ let $U_{\epsilon}:=\varphi\big{(}B_{d}(\mathbf{0},\epsilon)\big{)}$ denote be the image of the open ball of radius $\epsilon$. Since the singular support of $\mathfrak{Q}^{\tau}$ is the diagonal, the computation of the principal symbol of $\mathfrak{Q}^{\tau}$ at $(x,\xi)$ may be localized to $U_{\epsilon}$, meaning that in local coordinates it is given by the leading order term of the asymptotic expansion for $\xi\rightarrow\infty$ of the integral
$\int_{B_{d}(\mathbf{0},\epsilon)}e^{\imath\,\langle\xi,\mathbf{y}\rangle}\,\mathfrak{Q}^{\tau}(\mathbf{0},\mathbf{y})\,\rho(\epsilon^{-1}\,\mathbf{y})\,\mathcal{V}(\mathbf{y})\,\mathrm{d}\mathbf{y},$
where $\rho$ is a suitable fixed cut-off function identically equal to $1$ near the origin, and $\mathcal{V}(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is the pull-back by $\varphi$ of the Riemannian density on $M$ (here we occasionally blend intrinsic and local coordinate notation). If we pull this back by the dilation $\rho_{\epsilon}:\mathbf{x}\in B_{d}(\mathbf{0},1)\mapsto\mathbf{y}:=\epsilon\,\mathbf{x}\in B_{d}(\mathbf{0},\epsilon)$, we obtain the symbol at $(0,\epsilon\,\xi)$ of the corresponding $\mathfrak{Q}^{\tau}$ referred to the metric
$(\varphi\circ\rho_{\epsilon})^{*}(\kappa)=\epsilon^{2}\,g_{\epsilon},\quad\text{where}\quad g_{\epsilon}:=g_{st}+\sum_{|I|\geq 1}\epsilon^{|I|}\,x^{I}\,g_{I}.$ (161)
Now $g_{\epsilon}$ is a real-analytic Riemannian metric on $B_{d}(\mathbf{0},1)$, it is defined for sufficiently small $\epsilon$ and depends real analytically on $\epsilon$; hence the same holds of the corresponding operators $\mathfrak{Q}_{g_{\epsilon}}^{\tau}$ and their principal symbols. By (157), (160), and (161) we conclude that
$\sigma\left(\mathfrak{Q}_{g_{\epsilon}}^{\tau}\right)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}\cdot\big{(}1+\epsilon\,F_{m}(\xi)\big{)},$
where $F_{m}$ is homogeneous of degree 0 in $\xi$.
On the other hand, in view of (159) on $U_{\epsilon}$ we have
$\sigma\left(\mathfrak{Q}^{\tau}_{(\varphi\circ\rho_{\epsilon})^{*}(\kappa)}\right)(m,\epsilon\,\xi)=\epsilon^{d-1}\,\sigma\left(\mathfrak{Q}^{\tau/\epsilon}_{g_{\epsilon}}\right)(x,\epsilon\,\xi)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}\cdot\big{(}1+\epsilon\,F_{m}(\xi)\big{)}.$
Since the result must be independent of $\epsilon$, we conclude that $F_{m}=0$.
Summing up, we conclude the following:
Lemma 81.
The principal symbol of $\mathfrak{Q}^{\tau}$ is
$\sigma(\mathfrak{Q}^{\tau})(m,\xi)=\left(\frac{\pi\,\tau}{\|\xi\|}\right)^{\frac{d-1}{2}}.$
Going over the arguments in §2, §3 and especially §4 of [Z-2014] or §5, §6 and especially §7 of [Z-2020] in light of Lemma 81, one obtains the following description of $U_{\mathbb{C}}(t+2\,\imath\,\tau)$ as a dynamical Toeplitz operator.
  1. 1.
@@
    then on $X^{\tau}$ there exist a smoothly varying pseudodifferential operator $Q^{\tau}_{t}$ of degree (d1)/2-(d-1)/2 and a smoothly varying operator $R^{\tau}_{t}$ with $\mathcal{C}^{\infty}$ kernel such that
@@
    $\gamma^{\tau}_{t,0}(x)=(\pi\,\tau)^{\frac{d-1}{2}}\cdot e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}\cdot\langle\sigma_{J}^{(x)},\sigma_{J_{t}}^{(x)}\rangle^{-1}$ (164)
@@
The arguments in the proof of Theorem 7 apply, except that the leading term of the amplitude has been multiplied by a factor $(\pi\,\tau)^{(d-1)/2}\,(u\,\tau)^{-(d-1)/2}$ and the unitary factor $e^{\imath\,\tilde{\theta}^{\tau}_{t}(x)}$ in (164) replaces $e^{\imath\,\theta^{\tau}_{t}(x)}$. We have used that the principal symbol of $D^{\tau}_{\sqrt{\rho}}$ along $\Sigma^{\tau}$ (or, equivalently, of $\mathfrak{D}^{\tau}_{\sqrt{\rho}}$) is
In view of the rescaling $u\mapsto\lambda\,u$, this entails a further factor $\lambda^{-(d-1)/2}$. Furthermore, at the critical point (101) of the phase the product $u\,\tau=1$. Thus to leading order we obtain an extra overall factor $(\lambda/\pi\,\tau)^{-(d-1)/2}$. ∎